105. A 1.3-kg block is pushed up against a stationary spring compressing it a distance of 4.2 cm. When the block is released, the spring pushes it away across a frictionless, horizontal surface. What is the speed of the block, given that the spring constant of the spring is 1400 N/m?
1
Expert's answer
2020-11-17T13:01:46-0500

After the block is pushed up against a stationary spring, the potential energy accumulated in the spring is:


"U = \dfrac{kx^2}{2}"

where "k = 1400N\/m" is the spring constant and "x = 4.2cm = 0.042m" is the distance of compressing. After the block was released, all this energy was converted to the kinetic energy of the block:


"T = \dfrac{mv^2}{2} = U"

where "m = 1.3kg" is the mass of the block. Substituting the expression for "U", obtian:


"\dfrac{mv^2}{2} = \dfrac{kx^2}{2}\\
\space \\
v = x\sqrt{\dfrac{k}{m}}\\
v = 0.042\sqrt{\dfrac{1400}{1.3}} \approx 1.38m/s"

Answer. 1.38 m/s.